On the shear and moment diagrams can you find the x values where they cross the x-axis. Thank you!!
Draw the shear and moment diagrams for the beam. 600 lb/ft
Answer #1

4.980098- 1815X 121=x 95% 250 TWO 3600 lb ft qr0081- 97006 - 187=- 0=% ,21=X + 270016 V 121 k it & 12 - Bx TWO V/97009 #solutV Support Reactions : - G {Mp=0 – Ay x12 + ({ x 600x6}(12+3x6) +(x 600x12 ) * ( 12 x 3) =0|| -12 Ay + 25200 +28800 = 0 (Ay =My = - {u. . - (10086). . x (Mx = -50 (3) at x=0, (Mc =0) at x=6 , MA= - 50x6² MA=-3600 lb-ft) 1. For segment AB (GSX 18) W-Location of zero shear force Vx = 0 = 7200- 900x+2542 on solving (C = 12) My =-{. W.6. (x – 4) + Ay (x-6) - Į (W=Wk) (7–666